OpenStudy (anonymous):

If (4, y) is an ordered pair of the function, then y =

undeadknight26 (undeadknight26):

what is the function?

OpenStudy (the_fizicx99):

f(x) = 4 ?

OpenStudy (anonymous):

The rule y= -x + 7

undeadknight26 (undeadknight26):

isn't just: y = -4 + 7 y = 3 your pair is 4,3